Subject # 1 2 3 Male Male Elite Pro Male Cyclist Body Mass (kg) 100 57...

  1. Subject #

    1

    2

    3

    Male

    Male

    Elite Pro Male Cyclist

    Body Mass (kg)

    100

    57

    70

    Absolute VO2 max (L/min)

    3.5

    3.6

    6.1

    Ventilatory Threshold (as % of VO2max)

    47

    65

    88

    Watts @ VO2 Max

    220

    330

    450

    Watts @ Threshold

    85

    215

    400

    Power to Wt. Ratio (Watts/Kg)

    0.85

    3.5

    5.7

    At their Ventilatory Threshold

  2. Calculate the MET max for each subject.
Answer #1

1 MET =3.5 ml of O2 consumed/kg/min

Person 1----

3.5 L of O2/min=3500 ml of O2/min

Per kg = 3500/100=35 ml of O2/kg/min

It is equal to 10 MET.

Person 2 ----

3600 ml/min

Per kg = 3600/57 = 63.16 ml/kg/min

In MET =63.16/3 =21.05 MET

Person 3----

6100 ml/min

Per kg =6100/70= 87.14 ml/kg/min

In MET= 87.14/3 = 29.05 MET